# ANAIS–112 three years data: a sensitive model independent negative test of the DAMA/LIBRA dark matter signal

### v1

@@ Abstract
Weakly interacting massive particles (WIMPs) are well-motivated candidates for dark matter. One signature of galactic WIMPs is the annual modulation expected in a detector’s interaction rate, which arises from Earth’s rotation around the Sun. Over two decades, the DAMA/LIBRA experiment has observed such modulation with 250 kg of NaI(Tl) scintillators, in accordance with WIMP expectations but inconsistent with the negative results of other experiments. The signal, however, depends on the target, so in order to either validate or refute the DAMA signal it is necessary to replicate the experiment using the same material. This is the goal of the ANAIS–112 experiment, currently underway since August 2017 with 112.5 kg of NaI(Tl). In this work, we present a reanalysis of three years of data employing an improved analysis chain to enhance the experimental sensitivity. The results presented here are consistent with the absence of modulation and inconsistent with DAMA’s observation at nearly 3 C.L., with the potential to reach a 5 level within 8 years of data collection. Additionally, we explore the impact of different scintillation quenching factors in the comparison between ANAIS–112 and DAMA/LIBRA.
@@ Introduction
Among the myriad candidates proposed to constitute the dark matter (DM), Weakly Interacting Massive Particles (WIMPs) have attracted significant attention due to both, their compelling theoretical motivation and potential detectability [3].
@@
around the galactic center, the Earth’s rotation around the Sun introduces a minor correction to its velocity relative to the halo.
@@
This function is truncated at the Milky Way escape speed, =544 km s-1[12].
comprises three primary components [12]:
@@
galactic coordinates is given by km s-1, with =238 km s-1, (2) the Sun’s peculiar motion, which is km s-1 and (3)
@@
of the order of – cm2 when interpreted as a WIMP with spin and isospin-independent coupling. Such a signal should have already been observed by other
@@
cannot be calculated but must be measured and they have shown to depend on energy.
@@
In this work, we present a reanalysis of the data collected over the first 3 years, totaling an exposure of 312.12 kgyr.
@@
Furthermore, the sensitivity of the experiment has improved as anticipated in [52], confirming our expectations of achieving
@@
Considering energy-independent quenching factors, the results of ANAIS are incompatible with the DAMA signal at 2 C.L.
@@
for the rejection of high-rate periods, we remove events arriving within 1 s from a muon interaction in the veto. This criterion serves to reject potential secondary
particles generated by a cosmic muon in the detector or its shielding, and also avoids the long luminescence associated
with the muon’s passage through the scintillator, which could otherwise lead to numerous low-energy false events.
@@
In the figure, we also show our Monte Carlo (MC) background model (a comprehensive review can be found in [47]). The model takes as input independent estimates of the different radioactive contaminations. The background in the ROI is dominated by radioactive contamination of the NaI(Tl) crystal, particularly by 210Pb out of equilibrium. This isotope is found in varying quantities in the different crystals, both in the bulk and in the surface, being higher for the earliest detectors constructed, D0 and D1, at a level of 3.15 mBq kg-1. After introducing improvements in the growth and purification process, the 210Pb level in detectors D2 to D8 decreased to values in the range of 0.7–1.8 mBq kg-1. Another common contaminant of NaI(Tl) is 40K, due to its chemical affinity. In the ANAIS crystals, it is present at levels of around 1 mBq kg-1 and is responsible of the peak at 3.2 keV visible in the spectra of Fig. 1. This energy is released by the de-excitation of the atomic K-shell following an electron capture (EC), when the high energy ray escapes. Sometimes, this can hit another crystal, producing a coincident event that allows us to tag the low energy deposition and use it to both estimate the amount of 40K in the crystal and select these low-energy events for calibration and efficiency calculation. The cosmogenic isotope 22Na is also present in the crystals and produces a background in the ROI of similar origin, but in this case the K-shell relaxation energy is at 0.87 keV, and is only marginally present in the region between 1 and 2 keV. 3H is another cosmogenic isotope that contributes significantly to the ROI. It is a pure beta emitter with an end-point at 18.591 keV. Other cosmogenic isotopes of tellurium and iodine contribute to the region of interest through L- and K-shell EC emissions [53]. Their half-lives are shorter (between 10 and 150 days) and are not relevant to the total background of the experiment, but they are important for explaining the evolution of the background, especially in the latest detectors arriving at Canfranc (D6, D7, and D8). The agreement between the data and the model is very good down to 3 keV. Below this energy, there appears to be a component that is not well explained by the model. It could be optical noise that escapes the event selection or some radioactive background contribution missed in the model.
@@
As demonstrated in [51], the fit is unbiased, and the results do not exhibit dependence on the time bin size for values between 5 and 30 days.
ANAIS–112 results are consistent with the null hypothesis, with p-values of 0.31 and 0.66 for [1–6] and [2–6] keV energy regions, respectively.
@@
= –3.1 3.7 counts keV-1 ton-1 d-1 and 0.7 3.7 counts keV-1 ton-1 d-1, respectively.
@@
correspond to a sensitivity of 2.8 for each energy region.
@@
following [48], conveniently updated to the effective
@@
In this case, the sensitivity decreases to 2 accordingly to the narrower energy window, resulting in correspondingly lower statistics.
@@ Methods
The digitization scale is optimized for the study of the low energy events, so the events above 500 keV are out of the digitizer dynamic range and pulse area energy estimator, , saturates. For this reason, the ANAIS–112 DAQ system [29] incorporates a second signal line conveniently attenuated to retain information on the energy released by every high energy event through the use of charge-to-digital converters (QDC).
@@
Since there are no external sources of high energy available for calibrating the ANAIS–112 high energy regime, calibration of events above 100 keV is conducted independently for each background run using several easily identifiable peaks present in the background data. For every detector and run, the number of peaks used for calibration ranges from 6 to 7, depending on their presence in the background spectrum. Among them are: 238.6 keV (212Pb), 295.2 keV (214Pb), 351.9 keV (214Pb), 609.3 keV (214Bi), 1120.3 keV (214Bi), 1460.8 keV (40K) and 1764.5 keV (214Bi). Each peak is fitted to Gaussian lineshapes. Eventually, the calibration is performed via a linear regression between the nominal energies of the peaks and the Gaussian means using a second-degree polynomial.
Fig. 7 shows the high-energy calibrated background spectrum for single-hit events adding all the ANAIS–112 detectors over the first three years of operation. The upper panel shows the residuals ((fit – nominal)/nominal) for the positions of the main peaks identified in the background, being all of them below 1%.
@@
The 22Na/40K low energy peaks are excellent for low energy calibration, but their low rate and the low efficiency for the detection of the coincidence to select them accurately requires the accumulation of background data over long periods to observe them properly. It is also worth noting that the 22Na peak is below the analysis threshold, and despite efficiently triggering at the photoelectron level each PMT, the requirement of coincidence between the two PMTs within the 200 ns window results in a non-negligible decrease in trigger efficiency below 1 keV. This efficiency was estimated through a Monte Carlo simulation in [29] and has been used to correct the nominal energy of the 22Na peak from 0.87 keV to 0.90 keV. Thus, we have accumulated low-energy coincident events of 22Na and 40K over the first five years of measurement for each detector, and each peak has been fitted to a Gaussian lineshape added to a first-degree polynomial to estimate its mean value. Eventually, the ROI calibration has been conducted via linear regression between the Gaussian means and the nominal energies (0.90 and 3.2 keV, respectively) using a linear function.
@@
The trigger rate in the ROI is dominated by non-bulk scintillation events. For this reason, the development of robust protocols for the selection of events corresponding to bulk scintillation in sodium iodide is mandatory. Initially, the selection criteria applied in ANAIS–112 were based on standard cuts on a few parameters [29], and even though they demonstrated effectiveness above 2 keV, they showed weaknesses in the region from 1 to 2 keV. In order to improve the rejection of noise events between 1 and 2 keV, a machine-learning technique based on a Boosted Decision Tree (BDT) has been implemented. The detailed description of the BDT performance in ANAIS–112 with the old low-energy calibration can be found in [52]. Because of the implementation of the new low-energy calibration, the BDT filtering method requires a new training procedure using the updated populations. As training populations for BDT, we combine the following: as signal events, scintillation events ranging from [1–2] keV inside the crystal bulk produced by neutron interactions from dedicated neutron calibrations with a 252Cf source located outside the ANAIS–112 shielding, which are predominantly associated with elastic nuclear recoils in that energy region [58]; and as noise events, those coming from a blank module similar to the ANAIS–112 modules, but without NaI(Tl) crystal. This choice of the training populations is robust, as it entirely excludes background events, and uses bulk events as signal. The training process results in a newly constructed parameter named BDT, which combining the information of 15 discriminating parameters maximizes the separation between the signal and noise populations used in the training. For the event selection, we define an energy-dependent BDT cut, retaining only those events that exceed it. This selection criterion has been fine-tuned for each detector and energy bin. The corresponding efficiency is estimated individually for each detector by using 252Cf neutron calibration events. The ratio of events that pass the signal selection to the total events determines the acceptance efficiency which, when multiplied by the trigger efficiency, constitutes the total detection efficiency. Total efficiency for event detection in all the ANAIS–112 modules as a function of energy is shown in Fig. 9. The acceptance efficiency derived from the BDT cut notably exceeds (around 30% in [1–2] keV) that of the previous ANAIS–112 filtering procedure. Moreover, the BDT method significantly diminishes the background level below 2 keV for all detectors compared to that obtained using the previous ANAIS–112 protocols. In particular, the integrated rate from 1 to 2 keV is 5.390.04 and 4.400.03 c/keV/kg/d for the ANAIS–112 filtering procedure and the BDT method, respectively, representing an 18% reduction in background.
@@
Radioactive backgrounds and dark matter are expected to produce a constant rate of events in the detector when considering short time intervals. High-trigger rate periods may be caused by, for example, electrical or mechanical disturbances. In this scenario, high-rate periods, statistically inconsistent with the average detection rate of the experiment can be safely discarded. In order to do so, we evaluated the daily rate of low-energy single-hit events (below 3 keV) passing the BDT cut for each detector (blue line in Fig. 10). All bins exceeding three standard deviations from the detector’s average rate are removed. The red line in Fig. 10 shows the event rate surviving the cut. The fraction of rejected live time after applying this filtering varies for each detector but, in any case, it is less than 1%.
Data and code availability
ANAIS–112 will release all of the data included in this article, as done before, as well as convenient scripts for reproducing all the figures and results. They will be available at the website of the ORIGINS Excellence Cluster: https://www.origins-cluster.de/odsl/dark-matter-data-center/available-datasets/anais.

### v2

@@ Abstract
Weakly interacting massive particles (WIMPs) are well-motivated candidates for dark matter. One signature of galactic WIMPs is the annual modulation expected in a detector’s interaction rate, which arises from Earth’s revolution around the Sun. Over two decades, the DAMA/LIBRA experiment has observed such modulation with 250 kg of NaI(Tl) scintillators, in accordance with WIMP expectations but inconsistent with the negative results of other experiments. The signal depends on the target material, so to validate or refute the DAMA result, the experiment must be replicated using the same material. This is the goal of the ANAIS–112 experiment, currently underway since August 2017 with 112.5 kg of NaI(Tl). In this work, we present a reanalysis of three years of data employing an improved analysis chain to enhance the experimental sensitivity. The results presented here are consistent with the absence of modulation and inconsistent with DAMA’s observation at nearly 3 confidence level, with the potential to reach a 5 level within 8 years from the beginning of the data collection. Additionally, we explore the impact of different scintillation quenching factors in the comparison between ANAIS–112 and DAMA/LIBRA.
@@ Introduction
Among the myriad candidates proposed to constitute the dark matter (DM), Weakly Interacting Massive Particles (WIMPs) have attracted significant attention due to both their compelling theoretical motivation and potential detectability [3].
@@
around the galactic center, the Earth’s revolution around the Sun introduces a minor correction to its velocity relative to the halo.
@@
is truncated at the Milky Way escape speed [12], =544 km s-1. comprises three primary components [12]:
@@
galactic coordinates is given by , with =238 km s-1, (2) the Sun’s peculiar motion, which is km s-1 and (3)
@@
of the order of – cm2 when interpreted as a WIMP with spin- and isospin-independent coupling. Such a signal should have already been observed by other
@@
cannot be calculated but must be measured and they have been shown to depend on energy.
@@
As it will be explained in the last section of this article, differences in quenching factors are the only systematic effect that could compromise a direct comparison of experiments using the same target. In addition to improve the understanding of the scintillation quenching factors in NaI(Tl), this systematic effect could be handled by strongly reducing the energy threshold well below that of DAMA/LIBRA. ANAIS+ is one of the R&D projects pursuing this goal, by replacing the PMTs by SiPMs and operating the detector at 100 K.
@@
In this work, we present a reanalysis of the data collected over the first 3 years, totaling an exposure of 312.53 kgyr.
@@
Furthermore, the sensitivity of the experiment has improved as anticipated in Ref. [52], confirming our expectations of achieving
@@
Considering energy-independent quenching factors, the results of ANAIS are incompatible with the DAMA signal at 3 C.L.
@@
for the rejection of high-rate periods, we remove events arriving within 1 s from a muon interaction in the veto.
Scintillation time constants as long as 300 ms have been observed for high-energy events in NaI(Tl) [53], therefore this criterion helps to prevent the DAQ from triggering numerous low-energy false events after a muon’s passage through the scintillator. Additionally, it also rejects potential secondary particles generated by a cosmic muon in the detector or its shielding.
@@
In the figure, we also show our Monte Carlo (MC) background model (a comprehensive review can be found in [47]). The model takes as input independent estimates of the different radioactive contaminations. The background in the ROI is dominated by radioactive contamination of the NaI(Tl) crystal, particularly by 210Pb out of equilibrium. This isotope is found in varying quantities in the different crystals, both in the bulk and in the surface, being higher for the earliest detectors constructed, D0 and D1, at a level of 3.15 mBq kg-1. After introducing improvements in the growth and purification process, the 210Pb level in detectors D2 to D8 decreased to values in the range of 0.7–1.8 mBq kg-1. Another common contaminant of NaI(Tl) is 40K, due to its chemical affinity. In the ANAIS crystals, it is present at levels of around 1 mBq kg-1 and is responsible of the peak at 3.2 keV visible in the spectra of Fig. 2. This energy is released by the de-excitation of the atomic K-shell following an electron capture (EC), when the high energy ray escapes. Sometimes, this can hit another crystal, producing a coincident event that allows us to tag the low energy deposition and use it to both estimate the amount of 40K in the crystal and select these low-energy events for calibration and efficiency calculation. The cosmogenic isotope 22Na is also present in the crystals and produces a background in the ROI of similar origin, but in this case the K-shell relaxation energy is at 0.87 keV, and is only marginally present in the region between 1 and 2 keV. 3H is another cosmogenic isotope that contributes significantly to the ROI. It is a pure beta emitter with an end-point at 18.591 keV. Other cosmogenic isotopes of tellurium and iodine contribute to the region of interest through L- and K-shell EC emissions [54]. Their half-lives are shorter (between 10 and 150 days) and are not relevant to the total background of the experiment, but they are important for explaining the evolution of the background, especially in the latest detectors arriving at Canfranc (D6, D7, and D8). The agreement between the data and the model is very good down to 3 keV. Below this energy, there appears to be a component that is not well explained by the model. It could be optical noise that escapes the event selection or some radioactive background contribution missed in the model. Present studies with a larger digitization window (8 microseconds instead of 1.2) point to the first explanation as the main cause: this energy region seems to be dominated by a population of events with a time scale not compatible with NaI(Tl) scintillation and asymmetric in the light sharing between both PMTs. Light emissions at the PMTs (scintillation, corona discharges, etc.) could be responsible for these events. For the analysis presented in this paper, it acts as a background component whose evolution remains constant in time.
@@
The possible presence of a bias in the fit was studied in Ref. [51] using a large set of Monte Carlo pseudoexperiments, sampled from the background models with no modulation or with the modulation amplitude observed by DAMA/LIBRA. In all cases, the bias was found to be compatible with zero or negligible. In the present analysis, only the experimental data have changed with respect to Ref. [51], so the bias study remains valid in this case. The results do not exhibit dependence on the time bin size for values between 5 and 30 days.
ANAIS–112 results are consistent with the null hypothesis, with p-values of 0.40 and 0.64 for [1–6] and [2–6] keV energy regions, respectively.
@@
= –1.3 3.7 counts keV-1 ton-1 d-1 and 3.1 3.7 counts keV-1 ton-1 d-1, respectively.
@@
For illustrative purposes, Fig. 5 shows the fit results after subtracting the background term from Eq. 7 in both the fitting functions and the data for the energy region [1–6] keV for the combined data of the 9 detectors. The modulation observed by DAMA/LIBRA is shown in green.
@@
correspond to a sensitivity of (2.80.3) in [1–6] keV and (2.80.2) in [2–6] keV, where the uncertainty corresponds to the 68% C.L. DAMA/LIBRA uncertainty.
@@
following Ref. [48], conveniently updated to the effective
@@
In this case, the sensitivity is 3 because, although the statistics is reduced as a consequence of the reduction of the integration window, the signal-to-background ratio increases correspondingly.
@@ Methods
The digitization scale is optimized for the study of the low energy events, so the events above 500 keV are out of the digitizer dynamic range and pulse area energy estimator, , saturates because the digitized pulses (negative) are truncated at the minimum voltage (-1 V). For this reason, the ANAIS–112 DAQ system [29] incorporates a second signal line conveniently attenuated to retain information on the energy released by every high energy event through the use of charge-to-digital converters (QDC).
@@
Since there are no external sources of high energy available for calibrating the ANAIS–112 high energy regime, calibration of events above 100 keV is conducted independently for each background run using several easily identifiable peaks present in the background data. For every detector and run, which in average lasts for two weeks, the number of peaks used for calibration ranges from 6 to 7, depending on their presence in the background spectrum. Among them are: 238.6 keV (212Pb), 295.2 keV (214Pb), 351.9 keV (214Pb), 609.3 keV (214Bi), 1120.3 keV (214Bi), 1460.8 keV (40K) and 1764.5 keV (214Bi). Each peak is fitted to Gaussian lineshapes. Eventually, the calibration is performed via a linear regression between the nominal energies of the peaks and the Gaussian means using a second-degree polynomial.
Fig. 9 shows the high-energy calibrated background spectrum for single-hit events adding all the ANAIS–112 detectors over the first three years of operation. Panel b) shows the residuals ((fit – nominal)/nominal) for the positions of the main peaks identified in the background, being all of them below 1%.
@@
The 22Na/40K low-energy peaks are excellent for low energy calibration, but their low rate and the low efficiency for the detection of the coincidence to select them accurately requires the accumulation of background data over long periods to observe them properly. It is also worth noting that the 22Na peak is below the analysis threshold, and despite efficiently triggering at the photoelectron level each PMT, the requirement of coincidence between the two PMTs within the 200 ns window results in a non-negligible decrease in trigger efficiency below 1 keV. This efficiency was estimated through a Monte Carlo simulation in [29] and has been used to correct the nominal energy of the 22Na peak from 0.87 keV to 0.90 keV. Thus, we have accumulated low-energy coincident events of 22Na and 40K over the first five years of measurement for each detector, and each peak has been fitted to a Gaussian lineshape added to a first-degree polynomial to estimate its mean value. Eventually, the ROI calibration has been conducted via linear regression between the Gaussian means and the nominal energies (0.90 and 3.2 keV, respectively) using a linear function.
@@
The trigger rate in the ROI is dominated by non-bulk scintillation events. For this reason, the development of robust protocols for the selection of events corresponding to bulk scintillation in sodium iodide is mandatory. Initially, the selection criteria applied in ANAIS–112 were based on standard cuts on a few parameters [29], and even though they demonstrated effectiveness above 2 keV, they showed weaknesses in the region from 1 to 2 keV. In order to improve the rejection of noise events between 1 and 2 keV, a machine-learning technique based on a Boosted Decision Tree (BDT) has been implemented. The detailed description of the BDT performance in ANAIS–112 with the old low-energy calibration can be found in [52]. Because of the implementation of the new low-energy calibration, the BDT filtering method requires a new training procedure using the updated populations. As training populations for BDT, we combine the following: as signal events, scintillation events ranging from [1–2] keV inside the crystal bulk produced by neutron interactions from dedicated neutron calibrations with a 252Cf source located outside the ANAIS–112 shielding, which are predominantly associated with elastic nuclear recoils in that energy region [59]; and as noise events, those coming from a blank module similar to the ANAIS–112 modules, but without NaI(Tl) crystal. This choice of the training populations is robust, as it entirely excludes background events, and uses bulk events as signal. The fact that we do not have pure scintillation populations can slightly bias the training, although the major effect is underestimating the cut efficiency. The training process results in a newly constructed parameter named BDT, which combining the information of 15 discriminating parameters maximizes the separation between the signal and noise populations used in the training. For the event selection, we define an energy-dependent BDT cut, retaining only those events that exceed it. This selection criterion has been fine-tuned for each detector and energy bin (see [52] for details). The corresponding efficiency is estimated individually for each detector by using 252Cf neutron calibration events. The ratio of events that pass the signal selection to the total events determines the acceptance efficiency which, when multiplied by the trigger efficiency, constitutes the total detection efficiency. Total efficiency for event detection in all the ANAIS–112 modules as a function of energy is shown in Fig. 1. The acceptance efficiency derived from the BDT cut notably exceeds (around 30% in [1–2] keV) that of the previous ANAIS–112 filtering procedure. Moreover, the BDT method significantly diminishes the background level below 2 keV for all detectors compared to that obtained using the previous ANAIS–112 protocols. In particular, the integrated rate from 1 to 2 keV is 5.390.04 and 4.400.03 counts keV-1 kg-1 d-1 for the ANAIS–112 filtering procedure and the BDT method, respectively, representing an 18% reduction in background.
@@
Radioactive backgrounds and dark matter are expected to produce a constant rate of events in the detector when considering short time intervals. High-trigger rate periods may be caused by, for example, electrical or mechanical disturbances. In this scenario, high-rate periods, statistically inconsistent with the average detection rate of the experiment can be safely discarded. In order to do so, we evaluated the daily rate of low-energy single-hit events (below 3 keV) passing the BDT cut for each detector (blue line in Fig. 11). All bins exceeding three standard deviations from the detector’s annually averaged rate are removed. The red line in Fig. 11 shows the event rate surviving the cut. The fraction of rejected live time after applying this filtering varies for each detector but, in any case, it is less than 1%.
Data availability
The data used for this manuscript are available at the website of the ORIGINS Excellence Cluster: https://www.origins-cluster.de/odsl/dark-matter-data-center/available-datasets/anais.
Code availability
The scripts for reproducing all the figures and results are available at the website of the ORIGINS Excellence Cluster: https://www.origins-cluster.de/odsl/dark-matter-data-center/available-datasets/anais.